A. Judge Yankee Distilling Company produces a number of alcohol-based solvents for industrial contamination prevention use. One of the solvents, Ripit, is produced through an extensive distilling process. Yankee Distilling has developed the following standards for Ripit.

Standard Quantity
or Hours
Standard Price
or Rate
Standard Cost
Direct materials 2.50 ounces $ 22.00 per ounce $ 55.00
Direct labor 0.90 hours $ 16.00 per hour 14.40
Variable manufacturing overhead 0.90 hours $ 2.00 per hour 1.80
Total standard cost per unit $ 71.20

For the month of August, Yankee Distilling recorded the following activity related to the production of Ripit:

  1. Materials purchased, 14,000 ounces at a cost of $289,800.
  2. There was no beginning inventory of materials; however, at the end of the month, 4,050 ounces of material remained in ending inventory.

  3. The company employs 26 lab technicians to work on the production of Ripit. During August, they each worked an average of 150 hours at an average pay rate of $15.00 per hour.

  4. Variable manufacturing overhead is assigned to Ripit on the basis of direct labor-hours. Variable manufacturing overhead costs during August totaled $5,000.

  5. During August, the company produced 3,900 units of Ripit.

Required:

1. For direct materials:

a. Compute the price and quantity variances.

b. The materials were purchased from a new supplier who is anxious to enter into a long-term purchase contract. Would you recommend that the company sign the contract?

2. For direct labor:

a. Compute the rate and efficiency variances.

b. In the past, the 26 technicians employed in the production of Ripit consisted of 6 senior technicians and 20 assistants. During August, the company experimented with fewer senior technicians and more assistants in order to reduce labor costs. Would you recommend that the new labor mix be continued?

3. Compute the variable overhead rate and efficiency variances.
Answer #1
Actual materials used 9950 =14000-4050
Actual labor hours 3900 =26*150
1
a
Materials price variance 18200 F =289800-(14000*22)
Materials quantity variance 4400 U =22*(9950-3900*2.5)
b
Yes, as materials variance is favorable
2
a
Labor rate variance 3900 F =3900*(15-16)
Labor efficiency variance 6240 U =16*(3900-3900*0.9)
b
No, as labor variance is unfavorable
3
Variable overhead rate variance 2800 F =5000-(3900*2)
Variable overhead efficiency variance 780 U =2*(3900-3900*0.9)